The Polar equation r = theta represent a conic
1
Expert's answer
2020-05-07T17:38:55-0400

There are three conical sections:

ellipse with eccentricity "e:0\le e\le 1"

parabola with eccentricity "e:e=1"

hyperbola with eccentricity"e:e=1"

Grarh "r=\theta" in polar coordinates


"r=\theta" forms a spiral

"r=\theta" is not a conical section.